OpenStudy (anonymous):

Simplify. Assume b, c ≠o. (bc^3)^-3

OpenStudy (anonymous):

= b^-3*c^-9 = 1 /( b^3c^9)

OpenStudy (alfie):

\[(bc^3)^{-3} = \frac{1}{(bc^3)^3} \rightarrow \frac{1}{b^3c^9}\] Now you explain yourself the different from zero, because it wouldn't make sense if it was zero since 1/0 is undefined.

OpenStudy (anonymous):

ok so 1/b^3c^9 is the answer?

OpenStudy (alfie):

yup.

OpenStudy (anonymous):

kk THX